A map and GPS application for a smartphone was tested for accuracy. The


experiment yielded 26 error out of the 74 trials. Find the 95% confidence interval


estimate for the proportion of errors

1
Expert's answer
2021-11-04T19:12:54-0400

"\hat p=\frac{26}{74}=0.3514."

"95\%CI=(0.3514-1.96\sqrt{\frac{0.3514(1-0.3514)}{74}},0.3514+1.96\sqrt{\frac{0.3514(1-0.3514)}{74}})="

"=(0.2426,0.4601)."